Write the equation of a quadratic whose roots are -3/7 and 2/5. Answer without fractions.
Answer by TimothyLamb(4379)

(x + 3/7)(x - 2/5) = 0
xx - (2/5)x + (3/7)x - 6/35 = 0
xx - (2*7/5*7)x + (3*5/5*7)x - 6/35 = 0
xx - (14/35)x + (15/35)x - 6/35 = 0
xx + (1/35)x - 6/35 = 0
---
answer:
xx + 0.0285714285714x - 0.171428571429 = 0
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php